OpenStudy (anonymous):

why is 1/a-12 1/a^2 the same as a-12overa^2

OpenStudy (anonymous):

because it used the reciprocal ..

OpenStudy (anonymous):

\[\frac{ 1 }{ a }-12\frac{ 1 }{a ^{2}}=\frac{ a-12 }{ a ^{2} }why?\]

OpenStudy (anonymous):

ahh .. wrong interpretation on the first .. wait ..

OpenStudy (anonymous):

if i make a=2 does it work for u?

OpenStudy (anonymous):

hm , for me .. the best thing to do is to get the lcd of the left equation and try to solve it .. will you do that?

OpenStudy (anonymous):

\[\frac{ 1 }{a ^{2} }-12\frac{ 1 }{a ^{2} }?\]

OpenStudy (anonymous):

yes .. aha! try to solve it ..

OpenStudy (anonymous):

or \[\frac{ a }{a ^{2} }\]

OpenStudy (anonymous):